Operation Summary for October 21, 2009 - October 23, 2009 |
Wednesday Oct. 21 The day shift began with Operations monitoring the Tevatron (TeV) store 7284 (149.26E30) and the Recycler stash (125.0E10), with the Antiproton Source stacking (28.9E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. During the day shift, Linac personnel flushed the H- Source water resistor; this work caused a short downtime. During the evening shift, Operations received permission to send beam to the Meson Test (MTest) area. TeV personnel conducted end-of-store studies. Operations terminated store 7284 at 8:30 PM. Operations put the TeV into HEP shot setup and established store 7285 at 10:30 PM with an average initial luminosity (AIL) of 274E30. Operators searched and secured the MT6-1 & 2 enclosures in preparation for beam operations. Thursday Oct. 22 The midnight shift began with Operations monitoring store 7285 (119.09E30) and the stash (83.2E10), with Pbar stacking (19.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, a Linac RF station (LRF4) tripped off. It reset. During the day shift, Switchyard personnel established beam to the MTest area and conducted studies. Pbar personnel made a 2.5-hour access into Pbar to repair an Accumulator RF station (ARF4). Operations terminated store 7285 at 1:15 PM. Operations put the TeV into shot setup at 1:28 PM, but the shot was delayed due to a MI SBD failure. Operations established store 7286 at 3:56 PM with an AIL of 281.94E30. During the evening shift, the Tevatron Electron Lens (TEL1) tripped off on a vacuum interlock; it reset. Starting at 9:30 PM, a NuMI power supply (E:V118) began tripping off and taking down the NuMI beam permit; operators reset the supply and the interlock after each trip. Friday Oct. 23 The midnight shift began with Operations monitoring store 7286 (109.68E30) and the stash (178.2E10), with Pbar stacking (22.5E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, E:V118 continued to trip off and be reset. At 1:08 AM, a Booster vacuum pump (IPRF17) tripped off and took two RF stations down (BRF7 & 8); everything reset. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. |
|
More Information Comments and Suggestions |
| last modified 10/26/2009 email Fermilab |
| Security, Privacy, Legal | |